# ChartMaster: Advancing Chart-to-Code Generation with Real-World Charts and Chart Similarity Reinforcement Learning

### v1

@@ 4.3 Comparison with SOTA
Table 1 compares our ChartMaster-7B with various MLLMs on multiple chart-to-code benchmarks. ChartMaster-7B achieves SOTA results among open-source 7B-scale models, even approaching the performance of GPT-4o. Notably, ChartMaster-7B consistently outperforms the baseline Qwen2.5-VL-7B across all metrics; for example, on the ChartMimic benchmark, it improves both low-level and high-level metrics by nearly 40 percentage points, highlighting the effectiveness of our approach. Furthermore, although our training dataset is sourced from the much larger Qwen2.5-VL-72B model—essentially a distillation-like setting—ChartMaster-7B still surpasses Qwen2.5-VL-72B on several benchmarks. These results demonstrate the strength of the ChartMaster framework in narrowing, and in some cases closing, the performance gap with significantly larger models.
@@
65.5 39.9 34.2 67.4 43.8 4.60 2.18
$\checkmark$ 91.1 73.7 73.3 80.3 59.3 5.34 2.36
$\checkmark$ $\checkmark$ 93.8 78.2 77.3 88.2 62.6 5.65 2.46
4.4 Ablation Study
Ablation study on ChartMaster. To assess the contribution of each component, we conduct an ablation study as summarized in Table 2. The base Qwen2.5-VL-7B model, without ReChartPrompt or ChartSimRL, demonstrates limited performance across benchmarks, revealing its restricted ability in both code generation and visual/semantic understanding. Introducing the ReChartPrompt-240K dataset leads to significant improvements in all metrics, demonstrating the high quality and effectiveness of ReChartPrompt-240K for chart-to-code generation. Further applying ChartSimRL on top of ReChartPrompt brings consistent gains, especially in GPT-based evaluation scores. These results demonstrate that ChartSimRL effectively guides the model to produce code that better matches the original charts in both semantic and visual aspects.
@@ 4.4 Ablation Study
91.1 73.7 73.3
$\checkmark$ 92.1 76.2 75.2
$\checkmark$ 92.1 77.7 76.0
$\checkmark$ $\checkmark$ 93.8 78.2 77.3
@@
Baseline - 91.1 73.7 73.3
Precision $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}|}$ 90.0 72.6 72.9
Recall $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}^{*}|}$ 90.6 74.7 74.8
F1 $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{(|\mathcal{A}_{i}|+|\mathcal{A}^{*}|)/2}$ 91.6 75.4 75.3
Jaccard $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}\cup\mathcal{A}^{*}|}$ 92.1 76.2 75.2
@@
Ablation study on ChartSimRL. ChartSimRL introduces a novel multimodal chart similarity reward, which goes beyond conventional text-based rewards by integrating both semantic and visual similarity between the candidate and original charts. To analyze the contribution of each component, we conduct ablation experiments as shown in Table 4. The results indicate that using either $R_{i}^{\mathrm{attr}}$ or $R_{i}^{\mathrm{vis}}$ alone improves all performance. Among them, the visual similarity reward leads to more pronounced gains, highlights the critical role of visual fidelity in chart-to-code generation. Furthermore, combining both rewards yields the best results, demonstrating the effectiveness of a multi-faceted reward design.
Ablation study on Attribute Similarity. We further evaluate the impact of different attribute similarity metrics—Precision, Recall, F1 score, and Jaccard similarity—on the ChartMimic benchmark (Table 4). Results show that optimizing solely for Precision can slightly decrease certain metrics, indicating that focusing only on attribute correctness without considering coverage may be suboptimal. In contrast, Recall, F1, and Jaccard similarity all lead to better performance, with Jaccard similarity achieving the highest overall scores.
Notably, while the F1 score balances Precision and Recall through the harmonic mean, Jaccard similarity measures the intersection over union between predicted and reference attribute sets, imposing a stricter penalty on missing or redundant attributes. This stricter overlap requirement allows Jaccard similarity to better capture the true semantic similarity between attribute sets, thus more effectively guiding the model’s learning process.
Ablation Study on Visual Similarity. Table 5 presents an ablation study of various visual similarity metrics on the ChartMimic benchmark. Standard metrics like MSE, SSIM wang2004image, and PSNR hore2010image, mainly assessing pixel-level or structural similarity, generally perform worse than advanced methods. In particular, SSIM underperforms significantly, indicating that pixel-based measures struggle to capture the complex visual nuances required for chart generation.
@@
In contrast, CNN-based metrics like AlexNet krizhevsky2012imagenet, VGG simonyan2014very and ResNet he2016deep that compare features in a learned representation space consistently outperform the baseline and pixel-level metrics across all evaluation criteria. ResNet-18 achieves the highest performance, highlighting the effectiveness of deep visual features.
@@
Qwen-2.5-VL-72B 91.7 77.5 74.7
For MLLM-based metrics, we use high-level similarity prompts from ChartMimic with Qwen-2.5-VL-72B to evaluate the similarity between generated and reference charts. These metrics further improve upon the baseline, especially in Execution Rate and Low-Level similarity. However, they still fall short of the best CNN-based metrics in High-Level similarity, suggesting that while MLLMs have strong semantic understanding, further optimization is needed for specialized visual tasks.
Figure 3: Comparison of dataset diversity and model performance of Qwen2.5-VL-7B after fine-tuning with Chart2Code-160K and ReChartPromp-160K datasets.
Comparison with Advanced Dataset. To thoroughly evaluate the diversity and quality of our dataset, we compare it with Chart2Code-160K. For fairness, 160K samples are randomly selected from our data to form the ReChartPromp-160K subset. Results are shown in Fig. 3. The left subfigure displays normalized counts of unique attributes—Numeric, Color, Text, and Layout—extracted from both datasets using tool $\mathcal{G}(\cdot)$. Higher counts indicate greater attribute diversity. ReChartPromp-160K exhibits substantially more unique attributes across all categories, especially in Text and Layout. While Chart2Code-160K relies on seed data, often resulting in repetitive attributes (see Appendix Fig. 5), ReChartPromp-160K draws from 30,071 diverse arXiv papers, yielding richer and less redundant attribute distributions. This greater diversity leads to tangible benefits: as shown in the right subfigure, models trained on ReChartPromp-160K consistently outperform those trained on Chart2Code-160K, demonstrating the importance of attribute diversity for robust and effective chart understanding and code generation.
@@ 4.5 Qualitative analysis
To further evaluate the effectiveness of our proposed approach, Fig. 4 presents qualitative comparisons of chart outputs generated by different models on the ChartMimic benchmark. The results reveal several key insights: (1) The baseline model can generate basic chart layouts, but often fails to reproduce fine-grained visual details. Consequently, its outputs frequently diverge from the reference charts. (2) ReChartPrompt-240K enhances chart-to-code generation accuracy. Fine-tuning the base model with our ReChartPrompt-240K dataset (“Base.+ReCha.”) makes the generated charts more accurate and diverse. This shows that using real-world chart-code pairs helps the model better adapt to different chart types and styles. However, there are still minor issues, like slight color or element position mismatches, indicating that standard supervised learning alone cannot achieve perfect visual consistence. (3) ChartSimRL further improves visual and semantic alignment. Incorporating ChartSimRL algorithm enables the model to more closely align its outputs with the ground truth. For example, the color usage in chart 1 and the arrow style in chart 2 better match the ground truth.
In summary, by combining ReChartPrompt and ChartSimRL, ChartMaster achieves visually accurate and semantically precise chart code, demonstrating strong generalization. This integrated approach also significantly narrows the gap with models like GPT-4o, highlighting ChartMaster’s practical effectiveness in chart-to-code generation.
@@ 5 Conclusion
In this paper, we propose ChartMaster, including a novel chart-to-code generation framework and a customized reinforcement learning algorithm. By introducing ReChartPrompt, we overcome the data homogeneity issues of previous works and construct a highly diverse ReChartPrompt-240K dataset. Our ChartSimRL algorithm integrates both semantic and visual similarity rewards, enabling the model to generate chart code that closely matches the original charts. Experimental results demonstrate that ChartMaster achieves performance competitive with GPT-4o on the chart-to-code generation task.
@@
Beyond its technical contributions, ChartMaster can facilitate automated scientific reporting and support data-driven decision-making across a wide range of domains. While our current framework focuses on common chart types and Python-based code, extending its capabilities to support a broader variety of chart formats and programming languages remains an important direction for future work.
Appendix A Appendix

### v2

@@ 4.1 Comparison with SOTA
We instantiate ChartMaster on the Qwen2.5-VL-7B backbone, resulting in the ChartMaster-7B model, and conduct comprehensive comparisons with a range of MLLMs. The detailed implementation and evaluation protocols are provided in the Appendix A. As shown in Table 1, ChartMaster-7B achieves state-of-the-art performance among open-source models at the 7B scale, showing competitive performance against GPT-4o. Notably, ChartMaster-7B consistently outperforms the baseline Qwen2.5-VL-7B across all metrics; for instance, in the ChartMimic benchmark, it improves both low-level and high-level metrics by about 40 percentage points. Furthermore, although our training dataset is derived from the larger Qwen2.5-VL-72B model—essentially a distillation-like setting—ChartMaster-7B still surpasses Qwen2.5-VL-72B on several benchmarks. These results convincingly demonstrate the effectiveness of the ChartMaster framework.
4.2 Ablation Study
Ablation study on ChartMaster. To assess the contribution of each component, we conduct an ablation study as summarized in Table 2. The base Qwen2.5-VL-7B model, without ReChartPrompt or ChartSimRL, demonstrates limited performance across benchmarks, revealing its restricted ability in both code generation and visual/semantic understanding. SFT with the ReChartPrompt-240K dataset leads to significant improvements in all metrics, demonstrating the high quality and effectiveness of ReChartPrompt-240K for chart-to-code generation. Further applying ChartSimRL on top of ReChartPrompt brings consistent gains, demonstrating its effectiveness in guiding the model to generate code that more accurately captures the original charts both semantically and visually. These findings confirm the complementary strengths of ReChartPrompt and ChartSimRL in enhancing ChartMaster’s overall performance.
@@ 4.2 Ablation Study
65.5 39.9 40.7 67.4 43.8 4.60 2.18
$\checkmark$ 91.1 73.7 80.9 80.3 59.3 5.34 2.36
$\checkmark$ $\checkmark$ 93.8 78.2 85.1 88.2 62.6 5.65 2.46
@@
91.1 73.7 80.9
$\checkmark$ 92.1 76.2 83.9
$\checkmark$ 92.1 77.7 84.3
$\checkmark$ $\checkmark$ 93.8 78.2 85.1
@@
- - 91.1 73.7 80.9
Precision $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}|}$ 90.0 72.6 79.0
Recall $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}^{*}|}$ 90.6 74.7 81.7
F1 $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{(|\mathcal{A}_{i}|+|\mathcal{A}^{*}|)/2}$ 91.6 75.4 84.5
Jaccard $\frac{|\mathcal{A}_{i}\cap\mathcal{A}^{*}|}{|\mathcal{A}_{i}\cup\mathcal{A}^{*}|}$ 92.1 76.2 83.9
@@
Ablation study on ChartSimRL. ChartSimRL introduces a novel multimodal chart similarity reward that combines both semantic similarity ($R_{i}^{\mathrm{attr}}$) and visual similarity ($R_{i}^{\mathrm{vis}}$) between the candidate and original charts. To dissect the contribution of each component, we conduct ablation experiments summarized in Table 4. The results show that employing either $R_{i}^{\mathrm{attr}}$ or $R_{i}^{\mathrm{vis}}$ alone consistently improves performance across all evaluated metrics. Notably, the visual similarity reward yields more substantial gains, underscoring the critical importance of preserving visual fidelity in chart-to-code generation. Moreover, combining both rewards achieves the best overall results, demonstrating the advantage of a multi-faceted reward design that simultaneously captures semantic and visual aspects.
Ablation study on Attribute Similarity. We adopt Jaccard similarity as a stringent metric for attribute similarity, whereby a candidate table achieves a perfect score only if its attribute set exactly matches that of the ground truth; even minor discrepancies incur penalties. To thoroughly assess the impact of different attribute similarity measures—Precision, Recall, F1 score, and Jaccard similarity—we conduct experiments on the ChartMimic benchmark, with results summarized in Table 4.
Our findings indicate that optimizing exclusively for Precision may lead to a slight decline in overall performance, as the model can achieve high Precision by predicting a limited subset of correct attributes while neglecting overall coverage. In contrast, Recall emphasizes coverage, which helps mitigate this issue and yields modest improvements. The F1 score, by harmoniously balancing Precision and Recall, further alleviates extreme biases and delivers enhanced overall performance. Notably, Jaccard similarity, measuring the intersection over union between predicted and reference attribute sets, enforces stricter penalties on both missing and redundant attributes. This higher overlap requirement enables Jaccard similarity to more faithfully capture the true semantic similarity between attribute sets, thereby resulting in the best overall performance.
Ablation Study on Visual Similarity. We use ResNet-18 (He et al., 2016) to extract features from charts to compute visual similarity. In fact, there are numerous methods to measure the similarity between two charts. To investigate the impact of different visual similarity metrics on model performance, we conduct an ablation study summarized in Table 5. Standard metrics such as MSE, SSIM (Wang et al., 2004), and PSNR (Hore & Ziou, 2010) primarily evaluate pixel-level or structural similarity (details in Appendix B). The table shows that these metrics generally perform worse than more advanced methods. Notably, SSIM exhibits a significant decline in performance, indicating that pixel-based measures struggle to capture the complex visual nuances necessary for effective chart-to-code generation.
@@
In contrast, CNN-based metrics like AlexNet (Krizhevsky et al., 2012), VGG (Simonyan & Zisserman, 2014), and ResNet (He et al., 2016), which compare features in a learned representation space, consistently outperform both the baseline and pixel-level metrics across all evaluation criteria. Among them, ResNet-18 achieves the highest performance, highlighting the effectiveness of deep visual features.
For MLLM-based metrics, we leverage the high-level similarity prompt from ChartMimic combined with Qwen-2.5-VL-72B to evaluate the similarity between generated and reference charts. These metrics show improvements over the baseline. However, they still fall short of the best CNN-based metrics, suggesting that although MLLMs possess strong semantic understanding, further optimization is required for specialized visual tasks such as chart-to-code generation.
@@
Qwen-2.5-VL-72B 91.7 77.5 83.9
Figure 3: Comparison of diversity and fine-tuning results between Chart2Code-160K and ReChartPrompt-160K datasets.
Comparison with Advanced Dataset. To comprehensively evaluate the diversity and quality of our dataset, we compare it with Chart2Code-160K (Zhao et al., 2025). For a fair comparison, we randomly sample 160K instances from our full dataset to construct the ReChartPromp-160K subset. Using the attribute extraction tool $\mathcal{G}(\cdot)$, we count unique chart attributes—including numerical values, colors, textual elements, and layouts—in both datasets. A higher number of unique attributes indicates greater attribute diversity. As shown in the left panel of Fig. 3, ReChartPromp-160K exhibits a substantially richer attribute distribution across all categories, notably in text and layout. This advantage stems primarily from Chart2Code-160K’s reliance on seed data sources, which results in repeated attribute patterns, whereas ReChartPromp-160K samples from distinct arXiv papers, ensuring broader coverage and less redundancy (see Appendix Fig. 5). This higher diversity brings clear benefits: as illustrated in the right panel, models fine-tuned on ReChartPromp-160K consistently outperform those trained on Chart2Code-160K, demonstrating the importance of attribute diversity for robust and effective chart understanding and code generation.
@@ 4.3 Qualitative analysis
Based on extensive experiments, we observe that ReChartPrompt generates charts with diverse and rich attributes, enabling the construction of a high-quality dataset that substantially enhances model performance. Building upon the distinctive features of the chart-to-code generation task, we propose the ChartSimRL algorithm, which further enhances the model’s capabilities. To comprehensively analyze the improvements brought by these contributions, we conduct a qualitative comparison of generated charts at different training stages on the ChartMimic benchmark (Fig. 4). Our key findings are summarized as follows: (1) The baseline model produces basic chart layouts but often fails to replicate fine-grained visual details, leading to noticeable discrepancies between generated outputs and reference charts. (2) Fine-tuning the base model on our ReChartPrompt-240K dataset (“Base.+ReCha.”) significantly improves chart-to-code generation accuracy. This improvement arises from the diverse, high-quality training data generated by conditioning on real-world chart prompts. Nonetheless, minor issues remain, such as slight mismatches in color or element positioning compared to the ground truth, indicating that supervised fine-tuning alone does not achieve perfect visual consistency. (3) Incorporating the ChartSimRL algorithm further improves both visual and semantic alignment. Notably, the model demonstrates enhanced color accuracy (as seen in the first column of Fig. 4) and more faithful reproduction of arrow styles in the second column, reflecting improved attention to key factual details. (4) ChartMaster competes favorably with GPT-4o. Notably, the ChartMaster-7B model can generate charts that more closely resemble the ground truth than those from GPT-4o, especially excelling in “mimicking” chart attributes. Additional generation results in Appendix Fig. 8 consistently support these conclusions.
@@ 5 Conclusion
In this paper, we propose ChartMaster, a novel chart-to-code generation framework paired with a tailored reinforcement learning algorithm. By introducing ReChartPrompt, we address data homogeneity issues in prior work and build a highly diverse ReChartPrompt-240K dataset. Our ChartSimRL algorithm combines semantic and visual similarity rewards, enabling the model to generate chart code that closely matches original visuals. Experiments show ChartMaster achieves performance on par with GPT-4o in chart-to-code tasks. We will open source all resources to foster community development and advance research in this area.
@@
Beyond its technical innovations, ChartMaster supports automated scientific reporting and empowers data-driven decision-making across a wide range of domains. While our current framework targets common chart types and Python-based code, expanding its scope to include a wider range of chart formats and programming languages is an exciting direction for future work.
Appendix A Implementation and Evaluation Details
During the collection of arXiv papers, we explicitly exclude papers that are used as benchmarks to avoid potential data leakage. In the data generation stage, we apply a greedy sampling strategy to filter chart data, retaining only images of 12 predefined chart types and discarding all others. Then, we randomly select an instruction from $\mathbf{P}_{\text{rechart}}$ to prompt the Qwen2.5-VL-72B (Bai et al., 2025) model to generate code via nucleus sampling, with a temperature of 0.1 and a top-p of 0.9.
For training, we use the Qwen2.5-VL-7B model (Bai et al., 2025) in two stages. In Stage 1, we perform SFT on the entire ReChartPrompt-240K dataset with a learning rate of $2\times 10^{-5}$, batch size 128, and a cosine annealing scheduler for one epoch; the resulting model is saved for Stage 2. In Stage 2, ChartSimRL training is conducted on 10% of the ReChartPrompt-240K dataset, using a smaller learning rate of $5\times 10^{-6}$ and generating $M=4$ candidate codes per sample. Candidate sampling uses temperature 1.0, top-p 1.0, and top-k 80 to encourage diversity. The batch size remains 128 (32 samples $\times$ 4 candidates each).
For evalution, we assess the model’s chart-to-code generation performance on multiple benchmarks. ChartMimic Direct Mimic Task (Yang et al., 2024a): This benchmark includes 600 chart images. GPT-4o scores (0–100) serve as high-level similarity metrics. Additionally, low-level F1 scores for text, layout, chart type, and color are computed from code execution for fine-grained analysis. Plot2Code Direct Asking (Wu et al., 2024): Metrics include code pass rate, text match rate, and a 10-point GPT-4V visual similarity score, jointly assessing code correctness and visual fidelity. ChartX Chart Redrawing Task (Xia et al., 2024): This benchmark uses GPT-4 (0–5 scale) to evaluate code-generated chart redrawings.
Appendix B Standard Metrics
We consider two RGB images: the original chart image $I_{i}\in\mathbb{R}^{H\times W\times 3}$ and the generated chart image $\hat{I}_{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the height and width of the images respectively (both images are resized to the same height and width before comparison), and $3$ corresponds to the RGB color channels. Below, we describe how to quantify the visual similarity between $I_{i}$ and $\hat{I}_{i}$ using metrics such as Mean Squared Error (MSE), Structural Similarity (SSIM) (Wang et al., 2004), and Peak Signal-to-Noise Ratio (PSNR) (Hore & Ziou, 2010).
B.1 Mean Squared Error
The Mean Squared Error (MSE) is defined as:
$\mathrm{MSE}(I_{i},\hat{I}_{i})=\frac{1}{H\times W\times 3}\sum_{h=1}^{H}\sum_{w=1}^{W}\sum_{c=1}^{3}\left(I_{i}(h,w,c)-\hat{I}_{i}(h,w,c)\right)^{2}$
This formula computes the average squared difference between the pixel values of the two images over all spatial locations and color channels. A smaller MSE indicates higher similarity between $I_{i}$ and $\hat{I}_{i}$.
To convert the MSE into a similarity score, we define the MSE-based similarity as:
$\mathrm{MSE\_Similarity}=\frac{1}{1+\mathrm{MSE}(I_{i},\hat{I}_{i})}\in(0,1]$
  • When $\mathrm{MSE}\to 0$, $\mathrm{MSE\_Similarity}\to 1$, indicating the images are almost identical.
  • When $\mathrm{MSE}\to\infty$, $\mathrm{MSE\_Similarity}\to 0$, indicating large differences between the images.
B.2 Structural Similarity
The Structural Similarity (SSIM) is a perceptual metric that quantifies the similarity between two images by comparing local patterns of pixel intensities. It is computed on local sliding windows centered at each pixel location. For each window, local statistics including mean, variance, and covariance are calculated to evaluate the similarity. The final SSIM value for each channel is obtained by averaging these local SSIM values over all spatial positions, and the overall SSIM between two RGB images is computed by averaging over the three color channels.
Formally, for each color channel $c\in\{R,G,B\}$, the SSIM is defined as:
$\mathrm{SSIM}_{c}(I_{i}^{c},\hat{I}_{i}^{c})=\frac{(2\mu_{I_{i}^{c}}\mu_{\hat{I}_{i}^{c}}+C_{1})(2\sigma_{I_{i}^{c}\hat{I}_{i}^{c}}+C_{2})}{(\mu_{I_{i}^{c}}^{2}+\mu_{\hat{I}_{i}^{c}}^{2}+C_{1})(\sigma_{I_{i}^{c}}^{2}+\sigma_{\hat{I}_{i}^{c}}^{2}+C_{2})}$
where
  • $\mu_{I_{i}^{c}}$ and $\mu_{\hat{I}_{i}^{c}}$ are the local means computed within the sliding window.
  • $\sigma_{I_{i}^{c}}^{2}$ and $\sigma_{\hat{I}_{i}^{c}}^{2}$ are the local variances.
  • $\sigma_{I_{i}^{c}\hat{I}_{i}^{c}}$ is the local covariance.
  • $C_{1}=(K_{1}L)^{2}$ and $C_{2}=(K_{2}L)^{2}$ are constants to stabilize the division, with default values $K_{1}=0.01$, $K_{2}=0.03$. $L$ is the dynamic range of the pixel values. For 8-bit grayscale images, $L=255$. In our implementation, all images are converted to np.float32 and normalized by dividing by 255, so the pixel values are in the range $[0,1]$. Therefore, $L=1.0$ is used for SSIM calculation.
The overall mean SSIM between the two RGB images is then calculated by averaging over all spatial positions $(x,y)$ in each channel and then over the three channels:
$\mathrm{SSIM}(I_{i},\hat{I}_{i})=\frac{1}{3}\sum_{c=1}^{3}\frac{1}{H\times W}\sum_{x=1}^{H}\sum_{y=1}^{W}\mathrm{SSIM}_{c}(I_{i}^{c}(x,y),\hat{I}_{i}^{c}(x,y))\in[0,1]$
  • When $\mathrm{SSIM}\to 1$, the images are structurally almost identical.
  • When $\mathrm{SSIM}\to 0$, there are significant structural differences between the images.
B.3 Peak Signal-to-Noise Ratio
Peak Signal-to-Noise Ratio (PSNR) is a widely used metric to measure the quality of reconstructed images compared to the original images. It is defined as:
$\mathrm{PSNR}(I_{i},\hat{I}_{i})=10\log_{10}\left(\frac{L^{2}}{\mathrm{MSE}(I_{i},\hat{I}_{i})}\right)$
where $L$ is the dynamic range of the pixel values. For normalized images in $[0,1]$, $L=1.0$.
In practical scenarios, PSNR values typically range in tens of decibels and can vary widely, which may cause instability during optimization. To mitigate this effect, we normalize the PSNR values within each rollout batch by dividing them by the maximum PSNR value in that batch:
$\mathrm{PSNR}_{\mathrm{norm}}(I_{i},\hat{I}_{i})=\frac{\mathrm{PSNR}(I_{i},\hat{I}_{i})}{\max_{\hat{I}_{j}\in\text{rollout batch}}\mathrm{PSNR}(I_{j},\hat{I}_{j})}\in(0,1]$
  • When $\mathrm{PSNR}_{\mathrm{norm}}\to 1$, the reconstructed image $\hat{I}_{i}$ is very similar to the original image $I_{i}$.
  • When $\mathrm{PSNR}_{\mathrm{norm}}\to 0$, there exist significant differences between the images.